Meeting Information

  • 1930 UT (1330 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News

  • 20 Hz synchronization tests look good. Will move forward tomorrow.
  • EVLA Guides: http://evlaguides.nrao.edu
  • Meeting Fridays at 0845 for discussion of HW-related items (and decisions on where to push to Mainsaver tracking).
  • Staff Travel:
    • Michael: 27 Apr - 4 May
    • Lorant: 9 May - 22 May
    • Joan: 13 May - 18 May
    • Joe: 19 - 24 May
  • Arriving:
    • Crystal Brogan (NRAO-CV); RSRO
    • Todd Hunter (NRAO-CV); RSRO

Commissioning review/reports

Notes for week of 05 Apr 2010 (Black: old but still relevant; Blue: new)

  • Meta summary

Mode OPT Scheduling Executor/Corr Calibration strategy Archive CASA AIPS Data Distribution Other Overall Status
    OST not ready (done manually) OPT selected limits not possible - output averaged SBs affected by zero visibilities - needs check on short pol'n runs - can't handle mixed configs (no X band with other than Ka < 32 GHz)  

  • Went to a new system for the status of the OSRO observing/priorities. See:
  • Priorities notes (Commissioning, OSRO, RSRO, Demo science) (Note: Not changed from last week due to some set backs in the system.).
    • Key testing:
      • Multi-configuration SBs
      • Push to wider bandwidths
      • Switched power implementation (Meeting this week with Michael/Steve)

OSRO Priority Testing Status

System Verification

  • Array/Receivers (Rick)
    • EA12 has joined the array.
  • WIDAR update (Michael)
    • Multi-config/'hands-off' observing status
    • Narrow band checkout status
    • AC vs BD OSRO 2 status

Band Checkouts (ECDATA-55)

  • P Band Status: %Red%Testing
    • Antennas with no P-band Rx: 7, 16, 22, 24 (no feed); 1, 4, 5 Rxs are not working.
  • L Band Status: Ready
    • Frequency labeling (ECDATA-9, ECDATA-10): Ok
    • Sky orientation (ECDATA-10): Ok
    • Gain stability (ECDATA-23): Ok; some bandpass phase evolution seen; follow-up on really blank sky.
    • Tuning: updated.
    • Antennas with no L-band: EA06, EA20, EA26
    • Integrating down (ECDATA-23):
      • Averaging in time is ok; averaging between subbands has problems.
      • L band (really) blank field (ECDATA-40)
    • L band, HI checkout (ECDATA-47)

  • S Band Status: Ready.
    • S band edge tests (similar result to C for tunings near 4 GHz; potentially some satellite downlink RFI; review of aliasing/out-of-band intermodulation is under review (Rick)
    • Frequency labeling (ECDATA-24): Ok
    • Sky orientation (ECDATA-18): Ok
    • Gain stability (ECDATA-24): Ok (no jumps seen)
    • Tuning (ECDATA-24): Reviewing default frequency based on birdies; recommendation to shift to multiple of 128 MHz .
    • Integrating down (ECDATA-18): Ok
    • Antennas with S band: EA08, EA09, EA15, EA22, EA24, EA27, EA28

  • C Status: Ready
    • Frequency labeling: Ok. Will get more information from Eric's analysis.
    • Sky orientation (ECDATA-35): Ok
    • Gain stability (ECDATA-35): Ok
    • Tuning: Low frequency tuning issue; stepped up for last AH1004.
    • Integrating down (ECDATA-25):
    • Antennas with no C-band: EA13

  • X Band Status: Ready
    • Frequency labeling : ECDATA-17): Ok
    • Sky orientation (ECDATA-17): Ok
    • Gain stability (ECDATA-26): Ok (no jumps)
    • Tuning: No issues
    • Integrating down (ECDATA-26): Ok
    • Antennas
    • RFI: none seen

  • Q Band Status: Ready.
    • Need Q band suite of tests.

Mode testing

  • Polarimetry Status:
    • C (ECDATA-12): Ready
    • L (ECDATA-42): Ready
    • Eric reviewing AH1004 data for AIPS
    • George will be pushing things into JIRA shortly; noted that he is seeing an East/West difference in the data, don't settle to the same value; good candidate for a more detailed discussion (e.g., ionospheric effects, etc).
  • Spectral Line
    • Bandpass stability: ECSV-35
    • Bandpass tests: ECSV-92
      • Further tests in the queue
  • Wideband L (AB1345 - ) Sanjay
  • Solar (Rick) status and path forward.

OSRO Proposal Testing (ECDATA-34)

  • Status: https://spreadsheets.google.com/ccc?key=0Al6JLDuYrDsBdG90aTZoVlE4WGoyOXhNRWtyOXgzT2c&hl=en
    • AK725 (Ott)
    • AS1013 (Chomiuk)
    • AB1347 (Miller-Jones)
    • AC974, AC975, AR708, AW769 runs (Carilli)
    • AC980 (Aravena)
    • AI139 (Pannella)
    • AR704 (McMullin)
    • AS1021, AS1012 (Claussen)
    • AS1039 (Rupen)
    • AT378 (Myers)
    • AU129 (ECDATA-106), AU130 (ECDATA-148) (van Moorsel)
    • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [Butler]: Scan labeling synchronization. Any new instances seen? None reported
      • Med Priority [Perley]: Issues integrating down when averaging by channel. Okay on blank fields.
      • Med Priority [TBD]: Subband 0 problems: Being reviewed on Wednesday morning at the WIDAR meeting. Butler to report.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: Occasional zero-amplitude visibilities (needs software fix - what level). UVLOD and FITLD can handle this in AIPS; flagdata clip is effective in CASA; still need to understand and fix at the source.
      • Low Priority [TBD]: Occasional pure-noise scans.
      • Low Priority [Rupen/Pokorny]: Delay clunking; software fix pending.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

Commissioning plans for the week

System Tests

  • [Michael]: Enable multi-configuration SBs (with no hand-editing)
  • [Joan/Michael/Bryan]: Enable Dynamic Scheduling (diagnostics/debugging)
  • [Rick/Michael]: RSRO Observing Modes
  • Area Commissioning Tests
    • Group Leads should post plans to their respective JIRA tickets.
    • When SBs are developed, notify: Joe, Joan, Michael for prioritization in week's scheduling
    • Current priorities are:
      • WIDAR commissioning (driven by Michael); [TCOR; TOSR0001]
      • End-to-end or "Hand's off" observing (driven by Bryan)
      • General System Verification (driven by Rick)
      • Area Commissioning
        • 1. TOSR
        • 2. TSPE
        • 3. THIG/TMID/TLOW
        • 4. TPOL
        • 5. TSCH
        • ---
        • TCAL: plan proposed; needs review and scheduling.
        • TPLA/TSOL will need vetting soon. Need planned set of observations.
        • TDEM is complete for summer school (may still want mosaic for continuum polarimetry); other demo science will be done as filler projects.
        • TRAP need plan; currently ToO being handled via scheduling officer.

OSRO project tests

RSRO

  • Observations
  • Residents
    • Manuel Aravena (POC: Chris Carilli); area(s): High frequency checkout (reference pointing tests)
    • Crystal Brogan (POC: Claire Chandler); area(s): High frequency calibration
    • Laura Chomiuk (POC: Miller Goss; area(s): Correlator checkout; Spectral line observing)
    • Todd Hunter (POC: Claire Chandler; area(S): High frequency checkout
    • James Miller-Jones (POC: Claire Chandler; area(s): Demo Science)
    • Jacqueline van Gorkom (POC: Miller Goss; area(s): Correlator checkout; Spectral line observing)

Demo Science : ECDATA-153

  • Both test data sets obtained (3C391, IRC+10216)
    • Fed back to Summer School for reduction (Juergen)
  • 3C391 (James)
  • IRC+10216 (Mark)

Actions

  • [Rupen]: Develop daily/post-maintenance day checklists for the system.
  • [McMullin]: Post Eric's AIPS Guide as template for reduction strategy. Working on how to format this; twiki not good for providing a polished look to users. Want to use mediawiki; discussions with CV ongoing, to be resolved today.
  • [McMullin]: Add ECSO Projects to JIRA/tracking.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/Dhawan]: Develop test to develop/demonstrate readiness for high frequency observing.
  • [Perley]: Develop set of frequency accuracy tests (X band is complete).
  • [Perley]: Perform a frequency sweep test at S band to provide to engineers for review (Q2).
  • [Benson]: Update when automated handling of multiple spectral windows is available (Q2). Clarified: John will organize a small group this week (likely necessary candidates are Eric, Michael, Rick) to resolve the near and long term plans for handling the automated export to UVFITS for heterogeneous spectral windows (e.g., auto split by spectral window, etc). Results to be presented next Monday. Bryan - will lead this.

AOB

  • ECSO/RSRO Proposal status/policies (w Claire)

Planning review (for Q1/Q2 2010):

Note: Beginnings of group pages from the organization page:

Followup deferred. Begin to look at the detailed needs for the tests being scoped out by the leads. Want to finalize the detailed observing/testing plans by the end of this month and have a view of the longer term commitments needed. Establish an overview of the obstacles to deploying to the community; in particular, understand all of the details of the end-to-end data flow/processing to make decisions on what is working versus what is usable by the community. Something like:

Mode OPT Scheduling Executor/Corr Calibration strategy Archive CASA AIPS Data Distribution Other Overall Status
Example limit of 1 s int ok OPT selected limits not possible no recommendations avail ok can't cook coffee weather table - transient issue no-go

  • Critical Commissioning
    • System Verification (Perley/Rupen): Need plan to conclude complex gain, bandpass stability tests at L-Q, reference pointing, gain curves, tipping scans.
    • Observing Band Checkout (principally calibration strategies, flagging)
    • Observing Mode Support (broad modes based on science goals)
    • Observing support (e2e capability)
      • OSRO Testing (Gustaaf); ECSV-18 - needs to be pushed into JIRA.
      • Flux Calibrator Models (Amy); ECSV-43
  • Advanced Science Support

    • Need:
      • [Butler] Planning: Planetary observing (OPT setup; correlator needs); ECSV-7, ECSV-15
      • [Chandler] Planning: Demo Science
      • [Frail] Planning: ToO scheduling heuristics
      • [Sjouwerman] Planning: OPT tests for primary frequencies; ECSV-6
      • [van Moorsel] Planning: Data distribution plan
      • [Sjouwerman] Planning: Archive functionality tests; ECSV-11
      • [Mioduszewski] Planning: Calibrator survey; ECSV-27
      • [Brisken] Planning: Phased array (VLBI, pulsars); ECSV-22, ECSV-23
      • [Golap] Planning CASA multi-thread/cluster testing.
      • [Walker] Planning VLBI setup
-- JosephMcMullin - 19 Apr 2010
Topic revision: r3 - 2010-04-21, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback